There are 27 drops of conducting fluid .each drop has radius 'r' and each of them is charged to the same potential V1 . They  are combined to form a bigger drop the potential of the bigger drop is V2.Find the ratio V2/V1
 

Grade:10

2 Answers

Vikas TU
4 years ago
If q  be the charge on each drop 
Charge on big drop Q = 27 q 
R= radius of bigger drop 
4/3 πR^3 = 27 * 4/3 *π * r^3 
R=3r
V1 = kq/r 
V2 = kQ/R 
 
V2/V1 = 9 
V= kq/r 
 
Khimraj
4 years ago
volume will be constant
so 4/3 πR3 = 27 * 4/3 πr3
R=3r
V2/V1 = q2*r/q1*R
 =9
…................................................................